Solve the equation for exact values: 2cos^2(x) - 2sin^2(x) + 2 = 0
Select one:
a. π
b. 0
c. π/2
d. 2π

Answer by Alan3354(69443)

2cos^2(x) - 2sin^2(x) + 2 = 0
cos^2(x) - sin^2(x) + 1 = 0
---
cos^2 + sin^2 = 1
----
cos^2(x) - sin^2(x) + cos^2(x) + sin^2(x) = 0
2cos^2(x) = 0
cos(x) = 0
x = pi/2 ± n*pi, n = 0,1,2,3...